# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

### v1

@@ 2.3 Uniqueness Results for Non-rotating Bodies
Multiply it by $4\pi r^{3}\sigma(r)$ and integrate over $\mathbb{R}$, we obtain $4\pi\int_{0}^{\infty}r^{3}\left(\frac{d}{dr}A^{\prime}(\sigma)\right)\sigma dr=4\pi\int_{0}^{\infty}m(r)\sigma(r)rdr$. Easy to see $\int_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y=\int_{\{|x|\geq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y$. Then we compute
$\frac{G(\sigma,\sigma)}{2}$ $=\frac{1}{2}\left(\int_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y+\int_{\{|x|\geq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y\right)$
$=\int_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}d^{3}xd^{3}y$
$=\int\frac{(x)\mathbf{1}_{\{|x|\leq|y|\}}}{|x-y|}d^{3}xd^{3}y$
$=\int\sigma(y)d^{3}y\int\frac{\sigma(x)\mathbf{1}_{\{|x|\leq|y|\}}}{|x-y|}d^{3}x$
$=\int\sigma(y)V_{\left\{\sigma\mathbf{1}_{\{|x|\leq|y|)}\right.}d^{3}y$
@@
And $\sigma$ vanishes at 0 and outside a large ball, so is $\sigma A^{\prime}(\sigma)-A(\sigma)$. Then we use spherical coordinate transformation and integrating by parts and obtain
@@
Collect the results above we obtain $E_{0}(\sigma)=\int A(\sigma)-\frac{G(\sigma,\sigma)}{2}=\int(4A\left(\sigma)-3\sigma A^{\prime}(\sigma)\right)$. ∎
@@
Although the differentiability of $\sigma$ is used in Lieb and Yau’s proof [13], we refine their proof as above and do not necessarily need the differentiability of $\sigma$. This is the reason we do not need to assume (F4) holds. If $P(s)$ satisfies (F1)(F2)(F3), then thanks to Theorem 2.6, a minimizer $\sigma$ (after translation) meets the conditions in Lemma 2.38, thus (40) holds true for $\sigma$.
@@
In particular, if $P(\sigma)$ satisfies the polytropic equations of state, then $\sigma A^{\prime}(\sigma)=\gamma A(\sigma)=\frac{K\gamma}{\gamma-1}\sigma^{\gamma}$. Thanks to Lemma 2.38, we have $E_{0}(\sigma)=(4-3\gamma)U(\sigma)=(4-3\gamma)\int A(\sigma),G(\sigma,\sigma)=(6\gamma-6)U(\sigma)$. Actually, inspired by Lieb and Yau [13, Remark after Lemma 6], we can prove $G(\sigma,\sigma)=(6\gamma-6)U(\sigma)$ for minimizers using the scaling relations similar to those in in Section 3. That is, (with an abuse of notations) let $\sigma_{\lambda}(x)=\lambda^{3}\sigma(\lambda x)$ we have $\int\sigma_{\lambda}=\int\sigma$ thus $E_{0}(\sigma)\leq E_{0}\left(\sigma_{\lambda}\right),\left.\frac{d}{d\lambda}E_{0}\left(\sigma_{\lambda}\right)\right|_{\lambda=1}=0$. Moreover, $E_{0}\left(\sigma_{\lambda}\right)=\lambda^{3\gamma-3}U(\sigma)-\frac{\lambda}{2}G(\sigma,\sigma)$, then we have $\left.\frac{d}{d\lambda}E_{0}\left(\sigma_{\lambda}\right)\right|_{\lambda=1}=(3\gamma-3)U(\sigma)-\frac{G(\rho,\rho)}{2}=0$, which is what we want.
In order to obtain uniqueness result for minimizers, Lieb and Yau [13, Lemma 11 and Remark following] use proof of contradiction. They said the only property of $A(s)$ used was the convexity of $g$ (also see [16, Section 3]), where $g(s)=4A(s)-3A^{\prime}(s)s$. However, a closer look reveals that the proof needs the strict inequality $g^{\prime\prime}(s)<0$ on some interval, whereas convexity only gives $g^{\prime\prime}(s)\leq 0$. Nevertheless, it turns out we indeed can show that the convexity of $g$, together with the specific structure of $A$, implies $g^{\prime\prime}<0$ holds true in a small interval, which is already enough to make a contradiction in Proposition 2.43.
@@
Since $P^{\prime}(s)=sA^{\prime\prime}(s)$, we know $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$ is equivalent to $P(s)\in C^{2}\left(\mathbb{R}^{+}\right)$. $f^{\prime}(s)=3s^{2}A^{\prime\prime}\left(s^{3}\right)=\frac{3P^{\prime}\left(s^{3}\right)}{s},f^{\prime\prime}(s)=\frac{9s^{2}P^{\prime\prime}\left(s^{3}\right)\cdot s-3P^{\prime}\left(s^{3}\right)}{s^{2}}$. $f(s)$ is convex, thus $f^{\prime\prime}(s)=\frac{9s^{2}P^{\prime\prime}\left(s^{3}\right)\cdot s-3P^{\prime}\left(s^{3}\right)}{s^{2}}\geq 0$. If the conclusion does not hold, then $\exists\delta_{0}>0$, such that $\forall s\in\left(0,\delta_{0}\right),f^{\prime\prime}(s)=0$, which implies $3tP^{\prime\prime}(t)-P^{\prime}(t)=0$ for all $t\in\left(0,\delta_{0}^{3}\right)$, then $P^{\prime}(t)=Ct^{\frac{1}{3}}$ for some $C$ when $t\in\left(0,\delta_{0}^{3}\right)$. Then $\lim\limits_{t\rightarrow 0}\frac{P^{\prime}(t)}{t^{\frac{1}{3}}}=C$. Since we already know $P(t)$ satisfies (F2), that is $\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=0$, then $P\rightarrow 0$ as $t\rightarrow 0$. Then we can apply L’Hôpital’s rule [20, 5.13. Theorem] and obtain $0=\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=\lim\limits_{t\rightarrow 0}\frac{3}{4}\cdot\frac{P^{\prime}(t)}{t^{\frac{1}{3}}}=\frac{3C}{4}$, which means $C=0$. Then $P^{\prime}(t)=Ct^{\frac{1}{3}}=0$ when $t\in\left(0,\delta_{0}^{3}\right)$, notice again by ((F2)) $\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=0$, hence $P(t)=0$. It contradicts the assumption (F1) that $P(t)$ is strictly increasing for $t>0$. Therefore, we know $\forall\delta>0,\exists s_{\delta}\in(0,\delta)$, such that $f^{\prime\prime}\left(s_{\delta}\right)>0$. ∎
@@
If $P(s)$ satisfies (F1)(F2)(F3)(F4), $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$, and $A^{\prime}\left(s^{3}\right)$ is convex, then the minimizer of $E_{0}(\rho)$ is unique up to translation. In particular, if $P(s)$ satisifies the polytropic equations of state with index $\gamma>\frac{4}{3}$, the minimizer of $E_{0}(\rho)$ is unique up to translation.
@@
Let $f(s)=A^{\prime}\left(s^{3}\right),g(s)=4A(s)-3A^{\prime}(s)s$, by direct computation we obtain $f^{\prime\prime}(s)=6sA^{\prime\prime}\left(s^{3}\right)+9A^{\prime\prime\prime}\left(s^{3}\right)s^{4},g^{\prime\prime}(s)=-2A^{\prime\prime}(s)-3sA^{\prime\prime\prime}(s)$ when $s>0$. (Notice $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$ guarantees those derivatives exist and are continuous.) When $s>0,f$ is convex then $f^{\prime\prime}\geq 0$ implies $g^{\prime\prime}\leq 0$, then $g$ is concave. Therefore, we have $\frac{g(a)-g(b)}{a-b}\left\{\begin{array}[]{l}\leq g^{\prime}(b),b<a\\ \geq g^{\prime}(b),b>a\end{array}\right$., therefore, $\forall a,b\geq 0$, we have $g(a)-g(b)\leq g^{\prime}(b)(a-b)$.
If $m=0$, then $\sigma=0$ is unique. If $m>0$, without loss of generality, we assume $m=$ 1. Given two minimizers $\sigma$ and $\widetilde{\sigma}$, thanks to Theorem 2.6, after translation, we can assume $\sigma$ and $\widetilde{\sigma}$ are continuous, spherically symmetric, centered at 0, differentiable when they are positive, and radially decreasing. Then we know $\sigma(0)>0,\widetilde{\sigma}(0)>0$, and $\nabla\sigma(0)=\nabla\widetilde{\sigma}(0)=0$ (due to being symmetric and differentiable at 0). If $\sigma(0)=\widetilde{\sigma}(0)$, due to (31) (37) we have $\Theta_{\sigma}(0)=\Theta_{\widetilde{\sigma}}(0),\dot{\Theta}_{\sigma}(0)=\dot{\Theta}_{\widetilde{\sigma}}(0)$. By the uniqueness theorem of ODE, $\Theta_{\sigma}=\Theta_{\widetilde{\sigma}}$ where they are positive. Due to (37) we obtain $\sigma=\widetilde{\sigma}$. If $\sigma(0)\neq\widetilde{\sigma}(0)$, without loss of generality, we assume $\sigma(0)>\widetilde{\sigma}(0)>0$. Let $R>0$ and $\widetilde{R}>0$ be the radii of their supports. By Lemma 2.36, $m(r)>\widetilde{m}(r)$ for all $0<r<\max\{R,\widetilde{R}\}$. Then $R\leq\widetilde{R}$, for otherwise $m=\widetilde{m}(\widetilde{R})<m(\widetilde{R})<m$. Then since $\sigma$ and $\widetilde{\sigma}$ are minimizers, thanks to Lemma 2.38, use spherical coordinate transformation then we have
@@
Let $\sigma_{m}$ be minimizer for mass $m$, then the mass $m$ can be viewed as an strictly increasing function of minimizers’ central density.
@@
Now we have the existence and uniqueness result for a global minimizer of $E_{0}(\rho)$ (corresponding to non-rotating single star case). In the next 2 remarks we discuss the relation between single star system and binary stars system, and explain a bit the reason why we cannot find a global energy minimizer for $E_{J}(\rho)$, where $E_{J}(\rho)$ is given in (10).
Remark 2.45.
By the definition of $E_{J}(\rho)$, we know $\forall\rho\in{R}\left(\mathbb{R}^{3}\right),E_{J}(\rho)\geq E_{0}(\rho)\geq e_{0}(1)=e_{0}$. But $e_{0}$ cannot be attained by $E_{J}(\rho)$ if $J>0$. This is because $E_{J}(\rho)=E_{0}(\rho)+T_{J}(\rho)$. If $E_{0}(\rho)=e_{0}$, Theorem 2.6 (v) makes $I(\rho)>0$ thus $T_{J}(\rho)>0$, $E_{J}(\rho)>e_{0}$; if $E_{0}(\rho)>e_{0}$, then of course $E_{J}(\rho)>e_{0}$.
Remark 2.46.
Theorem 2.6 (ii) and Proposition 2.43 can help to construct $\rho\in R_{0}\left({R}^{3}\right)$ such that $E_{J}(\rho)-e_{0}<\epsilon$ where $\epsilon$ is arbitrarily small, and the test function can be chosen as $\rho(x)=\sigma_{m}(x)+\sigma_{1-m}(x-y)$ with $|y|$ sufficient large. The idea comes from Lieb, Morgan and McCann [12, 15, 17, 16].
Now we know the minimizer satisfies (EL) and it is unique. A natural question is: given a radial solution $\sigma$ to (EL) with central density (initial data) $\sigma(0)=\alpha$, is it a minimizer? It turns out the answer is yes, but we first need to discuss more about Remark 2.35 to tie together minimizers, radial solutions to (EL) and positive radial solutions to (38). We first review a lemma in [13] (but for simplicity, here we assume the pressure satisfies the polytropic equations of state).
Lemma 2.47 (Compact Support of $\Theta$ [13, Lemma 8]).
Assume the pressure satisfies the polytropic equations of state with index $\gamma>\frac{3}{2}$. Let $\Theta$ be the radial solution(i.e. $\Theta$ is spherically symmetric) to (38) obtained by integrating outwards from $r=$ 0 with the initial conditions $\Theta(0)=\beta>0$ and $\dot{\Theta}(0)=0$. Then there is some $R(\beta)\in(0,\infty)$ such that $\Theta(R(\beta))=0$. For $r<R(\beta),\dot{\Theta}(r)<0$. In particular, $\Theta(r)$ is decreasing up to $R(\beta)$.
@@
$\Theta(0)=\beta>0$ implies $R(\beta)\neq 0$ if $R(\beta)$ exists. Set $\rho=\left(A^{\prime}\right)^{-1}(\Theta)$, by Shell theorem, or the arguments similar as (27) and (29) we have $\Theta^{\prime}(r)=-4\pi\int_{0}^{r}\frac{t^{2}}{r^{2}}\rho(t)dt$, from which we get $\Theta^{\prime}(r)<0$ for $r<R(\beta)$. Integrate it again and use Fubini’s theorem we have
@@
Suppose $R(\beta)=\infty$, set $g(t)=t\rho(t)$, we could obtain the result that $g\in L^{1}([0,\infty),dr)$ [13, Lemma 8]. However, one can also show $\Theta(r)\geq\frac{c}{r}$ for large $r$ [13, Lemma 8]. Then $\rho=\left(A^{\prime}\right)^{-1}(\Theta)=C\Theta^{\frac{1}{\gamma-1}}\geq cr^{-\frac{1}{\gamma-1}}$ for large $r$. $\gamma>\frac{3}{2}$ implies $1-\frac{1}{\gamma-1}>-1$, thus $\int_{0}^{\infty}g(r)dr=\int_{0}^{\infty}r\rho(r)dr=\infty$, which make a contradiction. For more details one can refer to [13, Lemma 8 and Proposition 9]. ∎
Remark 2.48.
Actually $\dot{\Theta}(0)=0$ can be induced by the fact $\Theta$ is a radial solution.
@@
Remark 2.50.
Due to Remark 2.35, we know given a radial solution $\sigma$ to (EL) with compact support, we can find a positive radial solution $\Theta_{\sigma}$ to (38). Then Lemma 2.47 and Remark 2.49 tell us a radial solution $\Theta$ to (38) with positive initial data is only positve up to $r=R(\beta)$. Let
@@
We claim any radial solution $\sigma$ to (EL) has compact support. In fact, due to (36) we know $\sigma$ is decreasing when it is positive. If its support is unbounded, then $\sigma(r)>0$ for all $r\geq 0$. Then we can construct a solution $\Theta_{\sigma}$ to (38) with $\sigma=\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma}\right)$ (one can also see (37)), which implies $\Theta_{\sigma}$ also has unbounded support. But it contracts Remark 2.49.
Therefore, we can actually remove the compact support assumption at the beginning in this remark that “a radial solution $\sigma$ to (EL) with compact support”, and show the solutions to (EL) indeed correspond one-to-one with the positive solutions to (38).
Moreover, from $\lambda=\Theta(0)-V_{\sigma_{\Theta}}(0)$ above, one can see the choice of $\lambda$ in (EL) is not arbitrary but depends on the choice of $\Theta(0)$, which in turn is determined by the central density $\sigma(0)$ (Recall $\sigma=\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma}\right)$).
@@
Proposition 2.51 (Equivalence between (EL) Solutions and Minimizers [13, Lemma 12]).
Let $\sigma^{\alpha}$ be the unique nonnegative radial solution of (EL) with central density $\sigma(0)=\alpha$, then $\sigma^{\alpha}$ is the unique minimizer for $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$, where $m=\int_{\mathbb{R}^{3}}\sigma^{\alpha}$. That is, all radial solutions of (EL) parametrized by their central density are in fact minimizers of $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$ for some mass $m$.

### v2

@@ 2.3 Uniqueness Results for Non-rotating Bodies
Multiply it by $4\pi r^{3}\sigma(r)$ and integrate over $\mathbb{R}$, we obtain
$4\pi\int_{0}^{\infty}r^{3}\left(\frac{d}{dr}A^{\prime}(\sigma)\right)\sigma dr=4\pi\int_{0}^{\infty}m(r)\sigma(r)rdr$
Easy to see $\iint_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy=\iint_{\{|x|\geq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy$. Then we compute
$\frac{G(\sigma,\sigma)}{2}$ $=\frac{1}{2}\left(\iint_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy+\iint_{\{|x|\geq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy\right)$
$=\iint_{\{|x|\leq|y|\}}\frac{\sigma(x)\sigma(y)}{|x-y|}\,dx\,dy$
$=\iint_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{\sigma(x)\sigma(y)\mathbf{1}_{\{|x|\leq|y|\}}}{|x-y|}\,dx\,dy$
$=\int_{\mathbb{R}^{3}}\sigma(y)\,dy\int_{\mathbb{R}^{3}}\frac{\sigma(x)\mathbf{1}_{\{|x|\leq|y|\}}}{|x-y|}\,dx$
$=\int_{\mathbb{R}^{3}}\sigma(y)V_{\left\{\sigma\mathbf{1}_{\{|\cdot|\leq|y|\}}\right\}}(y)\,dy$
@@
And $\sigma$ vanishes at 0 and outside a large ball, so is $\sigma A^{\prime}(\sigma)-A(\sigma)$. Then we use spherical coordinate transformation and integration by parts, and obtain
@@
Collect the results above we obtain $E_{0}(\sigma)=\int_{\mathbb{R}^{3}}A(\sigma)\,dx-\frac{G(\sigma,\sigma)}{2}=\int_{\mathbb{R}^{3}}(4A\left(\sigma)-3\sigma A^{\prime}(\sigma)\right)\,dx$. ∎
@@
Although the differentiability of $\sigma$ is used in Lieb and Yau’s proof [42], we refine their proof as above and do not necessarily need the differentiability of $\sigma$. This is the reason we do not need to assume (F4) holds. If $P(s)$ satisfies (F1)(F2)(F3), then thanks to Theorem 2.6, a minimizer $\sigma_{m}$ (after translation) meets the conditions in Lemma 2.37, thus (39) holds true for $\sigma_{m}$.
@@
In particular, if $P(\sigma)$ satisfies the polytropic equations of state, then $\sigma A^{\prime}(\sigma)=\gamma A(\sigma)=\frac{K\gamma}{\gamma-1}\sigma^{\gamma}$. Thanks to Lemma 2.37, we have $E_{0}(\sigma)=(4-3\gamma)U(\sigma)=(4-3\gamma)\int_{\mathbb{R}^{3}}A(\sigma)\,dx$, and $G(\sigma,\sigma)=(6\gamma-6)U(\sigma)$. Actually, inspired by Lieb and Yau [42, Remark after Lemma 6], we claim
$G(\sigma,\sigma)=(6\gamma-6)U(\sigma)$
for minimizers using the scaling relations similar to those in Section 3. That is, (with an abuse of notations) let $\sigma_{\lambda}(x)=\lambda^{3}\sigma(\lambda x)$ we have $\int_{\mathbb{R}^{3}}\sigma_{\lambda}\,dx=\int_{\mathbb{R}^{3}}\sigma\,dx$ thus $E_{0}(\sigma)\leq E_{0}\left(\sigma_{\lambda}\right)$, $\left.\frac{d}{d\lambda}E_{0}\left(\sigma_{\lambda}\right)\right|_{\lambda=1}=0$. Moreover, $E_{0}\left(\sigma_{\lambda}\right)=\lambda^{3\gamma-3}U(\sigma)-\frac{\lambda}{2}G(\sigma,\sigma)$, then we have $\left.\frac{d}{d\lambda}E_{0}\left(\sigma_{\lambda}\right)\right|_{\lambda=1}=(3\gamma-3)U(\sigma)-\frac{G(\rho,\rho)}{2}=0$, and hence the claim follows.
In order to obtain uniqueness result for minimizers, Lieb and Yau [42, Lemma 11 and Remark following] use a proof by contradiction. They said the only property of $A(s)$ used was the convexity of $g$ (also see [46, Section 3]), where $g(s)=4A(s)-3A^{\prime}(s)s$. However, a closer look reveals that the proof needs the strict inequality $g^{\prime\prime}(s)<0$ on some interval, whereas convexity only gives $g^{\prime\prime}(s)\leq 0$. Nevertheless, it turns out we indeed can show that the convexity of $g$, together with the specific structure of $A$, implies $g^{\prime\prime}<0$ holds true in a small interval, which is already enough to make a contradiction in Proposition 2.42.
@@
Since $P^{\prime}(s)=sA^{\prime\prime}(s)$, we know $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$ is equivalent to $P(s)\in C^{2}\left(\mathbb{R}^{+}\right)$. Notice $f^{\prime}(s)=3s^{2}A^{\prime\prime}\left(s^{3}\right)=\frac{3P^{\prime}\left(s^{3}\right)}{s}$, and $f^{\prime\prime}(s)=\frac{9s^{2}P^{\prime\prime}\left(s^{3}\right)\cdot s-3P^{\prime}\left(s^{3}\right)}{s^{2}}$. Since $f(s)$ is convex, $f^{\prime\prime}(s)=\frac{9s^{2}P^{\prime\prime}\left(s^{3}\right)\cdot s-3P^{\prime}\left(s^{3}\right)}{s^{2}}\geq 0$. If the conclusion does not hold, then $\exists\delta_{0}>0$, such that $\forall s\in\left(0,\delta_{0}\right),f^{\prime\prime}(s)=0$, which implies $3tP^{\prime\prime}(t)-P^{\prime}(t)=0$ for all $t\in\left(0,\delta_{0}^{3}\right)$, then $P^{\prime}(t)=Ct^{\frac{1}{3}}$ for some $C$ when $t\in\left(0,\delta_{0}^{3}\right)$. Then $\lim\limits_{t\rightarrow 0}\frac{P^{\prime}(t)}{t^{\frac{1}{3}}}=C$. Since we already know $P(t)$ satisfies (F2), that is $\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=0$, then $P\rightarrow 0$ as $t\rightarrow 0$. Then we can apply L’Hôpital’s rule [52, Theorem 5.13] and obtain
$0=\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=\lim\limits_{t\rightarrow 0}\frac{3}{4}\cdot\frac{P^{\prime}(t)}{t^{\frac{1}{3}}}=\frac{3C}{4}$
which means $C=0$. Then $P^{\prime}(t)=Ct^{\frac{1}{3}}=0$ when $t\in\left(0,\delta_{0}^{3}\right)$, notice again by ((F2)) $\lim\limits_{t\rightarrow 0}\frac{P(t)}{t^{\frac{4}{3}}}=0$, hence $P(t)=0$. It contradicts the assumption (F1) that $P(t)$ is strictly increasing for $t>0$. Therefore, we know $\forall\delta>0,\exists s_{\delta}\in(0,\delta)$, such that $f^{\prime\prime}\left(s_{\delta}\right)>0$. ∎
@@
If $P(s)$ satisfies (F1)(F2)(F3)(F4), $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$, and $A^{\prime}\left(s^{3}\right)$ is convex, then the minimizer of $E_{0}(\rho)$ is unique up to translation. In particular, if $P(s)$ satisfies the polytropic equations of state with index $\gamma>\frac{4}{3}$, the minimizer of $E_{0}(\rho)$ is unique up to translation.
@@
Let $f(s)=A^{\prime}\left(s^{3}\right)$, and $g(s)=4A(s)-3A^{\prime}(s)s$, by direct computation we obtain $f^{\prime\prime}(s)=6sA^{\prime\prime}\left(s^{3}\right)+9A^{\prime\prime\prime}\left(s^{3}\right)s^{4}$, and $g^{\prime\prime}(s)=-2A^{\prime\prime}(s)-3sA^{\prime\prime\prime}(s)$ when $s>0$. (Notice $A(s)\in C^{3}\left(\mathbb{R}^{+}\right)$ guarantees those derivatives exist and are continuous.) When $s>0$, $f$ is convex then $f^{\prime\prime}\geq 0$ implies $g^{\prime\prime}\leq 0$, then $g$ is concave. Therefore, we have $\frac{g(a)-g(b)}{a-b}\left\{\begin{array}[]{l}\leq g^{\prime}(b),b<a\\ \geq g^{\prime}(b),b>a\end{array}\right$., therefore, $\forall a,b\geq 0$, we have $g(a)-g(b)\leq g^{\prime}(b)(a-b)$.
If $m=0$, then minimizer $\sigma_{m}=0$ is unique. If $m>0$, without loss of generality, we assume $m=$ 1. Given two minimizers $\sigma$ and $\widetilde{\sigma}$, thanks to Theorem 2.6, after translation, we can assume $\sigma$ and $\widetilde{\sigma}$ are continuous, spherically symmetric, centered at 0, differentiable when they are positive, and radially decreasing. Then we know $\sigma(0)>0,\widetilde{\sigma}(0)>0$, and $\nabla\sigma(0)=\nabla\widetilde{\sigma}(0)=0$ (due to being symmetric and differentiable at 0). If $\sigma(0)=\widetilde{\sigma}(0)$, due to (30) (36) we have $\Theta_{\sigma}(0)=\Theta_{\widetilde{\sigma}}(0)$, and $\dot{\Theta}_{\sigma}(0)=\dot{\Theta}_{\widetilde{\sigma}}(0)$. By the uniqueness theorem of ODE, $\Theta_{\sigma}=\Theta_{\widetilde{\sigma}}$ where they are positive. Due to (36) we obtain $\sigma=\widetilde{\sigma}$. If $\sigma(0)\neq\widetilde{\sigma}(0)$, without loss of generality, we assume $\sigma(0)>\widetilde{\sigma}(0)>0$. Let $R>0$ and $\widetilde{R}>0$ be the radii of their supports. By Lemma 2.35, $m(r)>\widetilde{m}(r)$ for all $0<r<\max\{R,\widetilde{R}\}$. Then $R\leq\widetilde{R}$; otherwise $m=\widetilde{m}(\widetilde{R})<m(\widetilde{R})<m$, which is a contradiction. Then since $\sigma$ and $\widetilde{\sigma}$ are minimizers, thanks to Lemma 2.37, use spherical coordinate transformation then we have
@@
Let $\sigma_{m}$ be the minimizer for mass $m$, then the mass $m$ can be viewed as a strictly increasing function of minimizers’ central density.
@@
Now we know the minimizer satisfies (EL) and it is unique. A natural question is: given a radial solution $\sigma_{(\alpha)}$ to (EL) with central density (initial data) $\sigma_{(\alpha)}(0)=\alpha$, is it a minimizer? It turns out the answer is yes, but we first need to discuss more about Remark 2.34 to tie together minimizers, radial solutions to (EL) and positive radial solutions to (37). We first review a lemma in [42] (but for simplicity, here we assume the pressure satisfies the polytropic equations of state).
Lemma 2.44 (Compact Support of $\Theta$ [42, Lemma 8]).
Assume the pressure satisfies the polytropic equations of state with index $\gamma>\frac{3}{2}$. Let $\Theta$ be the radial solution(i.e., $\Theta$ is spherically symmetric) to (37) obtained by integrating outwards from $r=$ 0 with the initial conditions $\Theta(0)=\beta>0$ and $\dot{\Theta}(0)=0$. Then there is some $R(\beta)\in(0,\infty)$ such that $\Theta(R(\beta))=0$. For $r<R(\beta),\dot{\Theta}(r)<0$. In particular, $\Theta(r)$ is decreasing up to $R(\beta)$.
@@
$\Theta(0)=\beta>0$ implies $R(\beta)\neq 0$ if $R(\beta)$ exists. Set $\rho=\left(A^{\prime}\right)^{-1}(\Theta)$, by Shell Theorem, or the arguments similar as (26) and (28) we have $\Theta^{\prime}(r)=-4\pi\int_{0}^{r}\frac{t^{2}}{r^{2}}\rho(t)dt$, from which we get $\Theta^{\prime}(r)<0$ for $r<R(\beta)$. Integrate it again and use Fubini’s theorem, we have
@@
Suppose $R(\beta)=\infty$, set $g(t)=t\rho(t)$, we could obtain the result that $g\in L^{1}([0,\infty),dr)$ [42, Lemma 8]. However, one can also show $\Theta(r)\geq\frac{c}{r}$ for large $r$ [42, Lemma 8]. Then for $r$ large enough we have
$\rho=\left(A^{\prime}\right)^{-1}(\Theta)=C\Theta^{\frac{1}{\gamma-1}}\geq cr^{-\frac{1}{\gamma-1}}$
Notice $\gamma>\frac{3}{2}$ implies $1-\frac{1}{\gamma-1}>-1$, thus $\int_{0}^{\infty}g(r)dr=\int_{0}^{\infty}r\rho(r)dr=\infty$, which make a contradiction. For more details one can refer to [42, Lemma 8 and Proposition 9]. ∎
@@
Due to Remark 2.34, we know given a radial solution $\sigma_{(\alpha)}$ to (EL) with compact support, we can find a positive radial solution $\Theta_{\sigma_{(\alpha)}}$ to (37). On the other hand, Lemma 2.44 and Remark 2.45 tell us a radial solution $\Theta$ to (37) with positive initial data is only positive up to $r=R(\beta)$. Let
@@
We claim any radial solution $\sigma_{(\alpha)}$ to (EL) has compact support. In fact, due to (35) we know $\sigma_{(\alpha)}$ is decreasing when it is positive. If its support is unbounded, then $\sigma_{(\alpha)}(r)>0$ for all $r\geq 0$. Then we can construct a solution $\Theta_{\sigma_{(\alpha)}}$ to (37) with $\sigma_{(\alpha)}=\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma_{(\alpha)}}\right)$ (one can also see (36)), which implies $\Theta_{\sigma_{(\alpha)}}$ also has unbounded support. But it contracts Remark 2.45.
Therefore, we can actually remove the compact support assumption at the beginning in this remark that “a radial solution $\sigma_{(\alpha)}$ to (EL) with compact support”, and show the solutions to (EL) indeed correspond one-to-one with the positive solutions to (37).
Moreover, from $\lambda=\Theta(0)-V_{\sigma_{\Theta}}(0)$ above, one can see the choice of $\lambda$ in (EL) is not arbitrary but is related to the choice of $\Theta(0)$, which in turn is related to the central density $\sigma_{(\alpha)}(0)=\alpha$ (Recall $\sigma_{(\alpha)}=\left(A^{\prime}\right)^{-1}\left(\Theta_{\sigma_{(\alpha)}}\right)$).
@@
Proposition 2.46 (Equivalence between (EL) Solutions and Minimizers [42, Lemma 12]).
Let $\sigma_{(\alpha)}$ be the unique nonnegative radial solution of (EL) with central density $\sigma(0)=\alpha$, then $\sigma_{(\alpha)}$ is the unique minimizer for $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$, where $m=\int_{\mathbb{R}^{3}}\sigma_{(\alpha)}\,dx$. That is, all radial solutions of (EL) parametrized by their central density are in fact minimizers of $E_{0}(\rho)$ among $mR(\mathbb{R}^{3})$ for some mass $m$.